OpenStudy (anonymous):

use logarithmic differentiation to find the derivative of the function y= sin^2 X * tan^4 X/(x^2+1)^2

OpenStudy (anonymous):

\(y= (sin^2 x) (tan^4\frac{x}{(x^2+1)^2}\)) Well, \(ln (ab)=ln a + ln b\) and also \(ln y^n=n ln y\) Have you attempted it?

OpenStudy (anonymous):

yes i have....thats not the equation .... y= (sin^2 x)(tan^4 x)/(x^2+1)^2

OpenStudy (anonymous):

Oh.That look much less complicated and easier. so \( ln y = 2 ln sin x + 4 ln tan x - ln(x^2 + 1)^2\) where \(\frac{d(ln y)}{dx}=\frac{y'}{y}\) So where were you stuck?

OpenStudy (anonymous):

i got....4(secx)^2 *(sinx)^2*(tanx)^3/(x^2+1)^2.....so i really didnt know how to incorparate log....so i took a difference route

OpenStudy (anonymous):

Ah. okay. As shown above, all you have to do is: example: \(\frac{d(ln sin x )}{dx}= \frac{cos x}{sin x} \)

OpenStudy (anonymous):

differentiate the term then put it on top of itself. You'll get something like this \(y' = y(2cot x + 8 csc 2x +\frac{4x}{x^2 + 1})\)